# A Comprehensive Evaluation of Large Language Models on Temporal Event Forecasting

### v1

@@
Abstract.
Recently, Large Language Models (LLMs) have demonstrated great potential in various data mining tasks, such as knowledge question answering, mathematical reasoning, and commonsense reasoning. However, the reasoning capability of LLMs on temporal event forecasting has been under-explored. To systematically investigate their abilities in temporal event forecasting, we conduct a comprehensive evaluation of LLM-based methods for temporal event forecasting. Due to the lack of a high-quality dataset that involves both graph and textual data, we first construct a benchmark dataset, named MidEast-TE-mini. Based on this dataset, we design a series of baseline methods, characterized by various input formats and retrieval augmented generation(RAG) modules. From extensive experiments, we find that directly integrating raw texts into the input of LLMs does not enhance zero-shot extrapolation performance. In contrast, incorporating raw texts in specific complex events and fine-tuning LLMs significantly improves performance. Moreover, enhanced with retrieval modules, LLM can effectively capture temporal relational patterns hidden in historical events. Meanwhile, issues such as popularity bias and the long-tail problem still persist in LLMs, particularly in the RAG-based method. These findings not only deepen our understanding of LLM-based event forecasting methods but also highlight several promising research directions. We consider that this comprehensive evaluation, along with the identified research opportunities, will significantly contribute to future research on temporal event forecasting through LLMs.
@@ Abstract.
ccs: Computing methodologies Temporal reasoningccs: Information systems Specialized information retrieval
1. Introduction
Figure 1. Illustration of leveraging LLM for temporal event forecasting. Given the complex event Israeli-Palestinian conflict , three formats of historical event representations, i.e., text (left side), graph (right side), or graph-text (both), can be fed into the LLMs, and the LLMs are expected to answer certain input questions about what will happend in the future.
@@ 1. Introduction
Temporal event forecasting aims to predict future events based on observed historical events (Zhao, 2022). This is a fascinating task that paves the way for humans to master the operating rules of the world, meanwhile, it also possesses significant practical and application value, such as offering early warning to critical events like civil unrest or regional conflicts (Leetaru and Schrodt, 2013; O’brien, 2010; Deng et al., 2020). Given the great values and impacts, temporal event forecasting has garnered increasing interest in both academic and research communities, and various studies have been conducted in recent years (Cai et al., 2022; Li et al., 2021b; Park et al., 2022; Ma et al., 2023b; Jin et al., 2021; Lv et al., 2020).
@@
Among these studies, the representative formulations of temporal events can be broadly categorized into three formats (Zhao, 2022): graph-based, text-based, and graph-text hybrid. Specifically, the graph-based approach represents each event, i.e., the so-called atomic event  (Dettmers et al., 2018; Shang et al., 2019; Schlichtkrull et al., 2018), in a structured format, i.e., a quadruple $(s,r,o,t)$, where $s$, $r$, $o$, and $t$ corresponds to the subject, relation  [^1] Relation refers to the type of the atomic event., object, and timestamp, respectively. As illustrated in Figure 1, each atomic event(e.g., (Palestine, use conventional military force, Israel, t-2)) is extracted from a textual document(e.g., Palestinian militant groups led by Hamas launch over 3,500 rockets from the Gaza Strip into Israel. ) , and multiple related atomic events form a complex event(Li et al., 2021c; Ma et al., 2023a, b), such as the Israeli-Palestinian conflict. Following such a structured representation, which is also termed as Temporal Knowledge Graph (TKG), various methods (Jin et al., 2020; Li et al., 2021b; Park et al., 2022; Ma et al., 2023b) that target at modeling the temporal and relational patterns have been proposed and achieved remarkable progress (Cai et al., 2022). In contrast, text-based methods, such as event script prediction (Lv et al., 2020) or forecast question answering (ForecastQA) (Jin et al., 2021), are characterized by directly consuming textual representations, which encapsulate more fine-grained details and contexts. Such detailed and contextual information is often ignored by graph-based methods since they are either not included in the ontology or not extracted by the information extraction system. Consequently, text-based event forecasting (Jin et al., 2021; Lv et al., 2020) concentrates more on the capabilities of natural language understanding and reasoning. The third branch, i.e., graph-text hybrid method, aims to take advantage of the merits from both formats: structured reasoning and fine-grained information. Nevertheless, existing graph-text hybrid methods (Deng et al., 2020, 2021; Ma et al., 2023a) primarily treat text as side information to be integrated into the graph-based backbones, without truly conducting reasoning and forecasting on the texts themselves.
@@
With the striking success of ChatGPT  [^2] https://chat.openai.com, LLMs have demonstrated impressive performance in various tasks (Kojima et al., 2022; Ouyang et al., 2022; Touvron et al., 2023) as well as event forecasting (Xu et al., 2023; Lee et al., 2023; Liao et al., 2023; Luo et al., 2024). These works pioneer in applying LLMs to the task of event forecasting through in-context-learning (ICL) (Lee et al., 2023), instruction tuning (Xu et al., 2023; Luo et al., 2024), and retrieval augmented generation (RAG) (Sun et al., 2023; Liao et al., 2023). However, on the one hand, most of these works only focus on the graph-based formulation (Liao et al., 2023; Luo et al., 2024) through discrete graph reasoning, while ignoring the broad contextual information retained by natural language. Thereby, integrating the raw texts, from which the structured events are extracted, into the LLM-based forecasting processing is a natural and rational direction. On the other hand, considering that event forecasting is mostly related to critical scenarios such as international relationship and domestic stability, trustability and explainability are essentially required, while the solutions that solely rely on the internal knowledge of LLMs would suffer from the problem of hallucination (Lee et al., 2023). Despite RAG (Lewis et al., 2020) is a promising solution, the current graph-based RAG (Liao et al., 2023; Sun et al., 2023) is prone to noisy events that are obtained by unreliable event extraction systems. Therefore, the performance of text-based or graph-text hybrid RAG on temporal event forecasting is worth of further exploration.
Moreover, several previous works, such as SeCoGD[27], have identified that event forecasting suffers from severe popularity bias. Specifically, conventional methods harness historical events to provide additional contextual information for event prediction. Even without relevant historical events, LLMs can still provide correct answers since they implicitly perceive extensive knowledge in their parameters. As shown in Figure 1, although Iraq did not appear in previous knowledge graphs, the model’s understanding of the relationship between Iraq and Iran led to the correct answer. However, when the object to be predicted is an entity that appears for the first time or rarely, such as Egypt in the figure, the performance of LLMs remains uncertain.
@@
To fill this gap, we aim to conduct a comprehensive study to evaluate the performance of LLMs on temporal event forecasting. Nevertheless, the main obstacle lies in the lack of well-established benchmark datasets. Most existing works are using graph-only datasets, e.g., GDELT (Leetaru and Schrodt, 2013) and ICEWS (O’brien, 2010), or text-only datasets, e.g., ForecastQA (Jin et al., 2021). There are some text-enriched event graph datasets (Ma et al., 2023a, b), which enrich an existing structured TKG dataset with the raw texts that are downloaded through the news article URLs offered by GDELT. Unfortunately, the structured events in these datasets are highly noisy, because they are either from the original GDELT dataset (Ma et al., 2023a) or extracted using poor-performing systems (Ma et al., 2023b), with only about 50% accuracy as demonstrated in Section 2.2.
@@
In this work, to conduct the evaluation, we first construct an exploratory benchmark data based on a previous dataset MidEast-TE (Ma et al., 2023b) and name our dataset as MidEast-TE-mini (short as MidEast-TE-m). Specifically, we sample a subset from the MidEast-TE (Ma et al., 2023b) and extract structured events using the state-of-the-art (SOTA) LLM GPT-4  [^3] https://platform.openai.com/docs/models/gpt-4-and-gpt-4-turbo. Based on this dataset, we design a series of baseline methods to evaluate the performance of LLMs on temporal event forecasting. We are particularly interested in the following research questions: (1) how do LLMs perform on different formulations of event forecasting, including different modalities of input as well as forecasting objectives? Additionally, how are different LLM backbones and fine-tuning affecting the forecasting results? (2) how does RAG perform on event forecasting and how do different retrieval models and settings affect the forecasting performance? (3) how do various methods perform w.r.t. popularity bias (long-tail) issues? To answer these research questions, we conduct extensive experiments to verify the functionality of every designed component. The experimental results unveil multiple interesting phenomena, which not only deepen our understanding of LLM-based event forecasting methods but also surface several interesting research directions. The main contributions of this work are as follows:
@@
  • To the best of our knowledge, we are the first to systematically evaluate LLM-based methods on text-involved temporal event forecasting.
@@ 2.1. Problem Formulation
Conventional temporal event forecasting methods define each event as a quadruple $(s,r,o,t)$, referred to as an atomic event. Following recent studies (Li et al., 2021c; Ma et al., 2023b), atomic events are grouped into complex events (CE), which include a set of correlated atomic events and represent major events that span over longer time spans and cover more entities. For example, all the events within the red dashed box in Figure 1 illustrate a complex event, specifically the Israeli-Palestinian conflict. Consequently, we define an atomic event at timestamp $t$ as $(s,r,o,t,c)$, where $s\in\mathcal{E}$, $r\in\mathcal{R}$, and $o\in\mathcal{E}$ represent the subject, relation, object, and timestamp respectively; $\mathcal{E}$ and $\mathcal{R}$ are the entity set and relation set, and $c$ denotes the complex event type. Notably, the timestamp $t$ is converted into a relative timestamp. All the atomic events at the same timestamp $t$ form an event graph denoted as $G_{t}=\{(s_{n},r_{n},o_{n},t,c_{n})\}_{n=1}^{N}$, where $(s_{n},r_{n},o_{n},t,c_{n})$ denotes the n-th event and $N$ is the number of events.
@@
In addition to the pure graph-based representation, we aims to incorporate raw news text into temporal event forecasting. Specifically, each event graph $G_{t}$ is extracted from a list of news documents $D_{t}=\left[d_{1},d_{2},\ldots,d_{k}\right]_{k=1}^{K}\in\mathcal{D}$, where $d_{k}$ denotes the k-th document at timestamp $t$ and $\mathcal{D}$ denotes the document set. It is noteworthy that the same document $d$ can be involved in multiple events. Given the historical event graphs or news documents prior to $t$, denoted as $\mathbf{G}_{<t}=\left\{G_{0},G_{1},\ldots,G_{t-1}\right\}$ and $\mathbf{D}_{<t}=\left\{D_{0},D_{1},\ldots,D_{t-1}\right\}$, respectively, and a query, represented as $(s,r,t)$, temporal event forecasting aims to predict the missing object.
2.2. Dataset Construction
@@
We first describe the data source that we utilized for curating the dataset. Then we present the dataset construction pipeline, followed by the human evaluation results.
2.2.1. Data Source
@@
We construct our dataset based on a subset of MidEast-TE(Ma et al., 2023b), therefore, we name our dataset as MidEast-TE-mini, short as MidEast-TE-m. The original MidEast-TE dataset includes both structured atomic events and news articles, which have already been clustered into complex events. Given the large scale of MidEast-TE, we sample a subset of complex events from it and keep the associated news articles as the original documents to build our dataset. Specifically, the time spans of the complex events in MidEast-TE are of high divergence, ranging from several days to more than three months. To reduce the potential bias introduced by outliers, we just sample 120 complex events whose time spans are within 40-60 days. For the shorter and longer complex events, we do not cover them in this exploratory dataset and leave them for future work.
2.2.2. Construction Pipeline
@@
Event Extraction. We instruct GPT-4 to perform the task of event extraction. Since the set of event types is quite large (¿200), putting all the event definitions into prompts yields high cost and may decrease the performance. Therefore, we follow the previous approach (Ma et al., 2023b) and conduct a multi-level event extraction given the three-layer hierarchical structure of CAMEO (Boschee et al., 2015) ontology. We include the description of each event type defined in CAMEO in the prompt. During the first-level extraction, we break down each article into multiple sentence-level trunks, each with around 150 tokens. Then, we extract the first-layer atomic events from each trunk. After recognizing atomic events based on first-level event types, for the second-level extraction, we group 15 atomic events with the same first-level event type and their respective trunks into a single prompt. This approach aims to maximally share the prompt while maintaining the high-quality extraction. The same procedure is then applied to the third-level extraction.
@@
2.2.3. Human Evaluation
Table 1. Statistics of our curated dataset MidEast-TE-m
@@ 2.2.3. Human Evaluation
In order to evaluate the quality of the benchmark dataset, we conduct human evaluation. Specifically, given a news article and the extracted events, we ask the human evaluator to tell whether the extracted events are valid or not based on the original article. We randomly sample 20 documents from the document set of the dataset, and conduct human evaluation based on the following criteria:
@@
To be noted, for every extracted atomic event, we evaluate it once by going through the three criteria sequentially. For example, if we identify the Time is incorrect, we stop the checking of the following two criteria. Table 2 shows the accuracy and error types of the event extraction results, including both our dataset MidEast-TE-m and previous datasets of MidEast-TE (Ma et al., 2023b) and GDELT-TE (Leetaru and Schrodt, 2013). To be noted, MidEast-TE is constructed using Vicuna-7b (Chiang et al., 2023), while GDELT-TE is curated using proprietary extraction system. Clearly, the dataset constructed by our pipeline is the most accurate one. In our dataset constructed using GPT-4, there are the fewest extraction errors, which makes the following forecasting research more valid and reliable. However, we admit that an overall accuracy of 72.6% is still unsatisfactory in practice, and further efforts will be devoted to improved event extraction performance.
Table 2. Error analysis of the event extraction results in different datasets.
Dataset #atomic events Acc.(%) error type (%) time relation entity GDELT-TE (Leetaru and Schrodt, 2013) 148 29.73 3.85 31.73 64.42 MidEast-TE (Ma et al., 2023b) 35 55.56 0 92.86 7.14 MidEast-TE-m(ours) 78 72.60 16.67 27.78 55.56
3. Methods
Figure 2. Illustration of rule-based history and retrieved history. The rule-based history is constructed by a set of predefined rules. In contrast, the retrieved history dynamically searches context from the temporal knowledge graph or news documents according to the current query.
Table 3. Input format of history
Graph-only (Pakistan Foreign Ministry, Cooperate economically, Egyptian counterpart officials, 2190); (Shah Mahmood Qureshi, Express intent to provide economic aid, Egypt FM Shoukry, 2190)
Text-only [Date]2190: Pakistani Foreign Minister Shah Mahmood Qureshi expressed great compatibility in visions between Cairo and Islamabad, highlighting the potential for cooperation in eradicating terrorism and intolerance during his visit to Egypt.
Mixed [Date]2189: Foreign Minister Shah Mahmood Qureshi visited Egypt for a two-day official visit to discuss bilateral relations and economic cooperation with his Egyptian counterpart, Sameh Hassan Shoukry, and to promote trade relations with Africa. (Shah Mahmood Qureshi, Make optimistic comment, Pakistan); (Egyptian Ambassador, Host a visit, Chief of Army Staff (COAS) General Qamar Javed Bajwa)

### v2

@@
Abstract
Recently, Large Language Models (LLMs) have demonstrated great potential in various data mining tasks, such as knowledge question answering, mathematical reasoning, and commonsense reasoning. However, the reasoning capability of LLMs on temporal event forecasting has been under-explored. To systematically investigate their abilities in temporal event forecasting, we conduct a comprehensive evaluation of LLM-based methods for temporal event forecasting. Due to the lack of a high-quality dataset that involves both graph and textual data, we first construct a benchmark dataset, named MidEast-TE-mini. Based on this dataset, we design a series of baseline methods, characterized by various input formats and retrieval augmented generation (RAG) modules. From extensive experiments, we find that directly integrating raw texts into the input of LLMs does not enhance zero-shot extrapolation performance. In contrast, fine-tuning LLMs with raw texts can significantly improve performance. Additionally, LLMs enhanced with retrieval modules can effectively capture temporal relational patterns hidden in historical events. However, issues such as popularity bias and the long-tail problem persist in LLMs, particularly in the retrieval-augmented generation (RAG) method. These findings not only deepen our understanding of LLM-based event forecasting methods but also highlight several promising research directions. We consider that this comprehensive evaluation, along with the identified research opportunities, will significantly contribute to future research on temporal event forecasting through LLMs.
keywords:
@@ Abstract
[…]
[1]organization=School of Computer and Cyber Sciences, Communication University of China, city=Beijing, postcode=100024, country=China
[…]
[2]organization=Department of Computer Science, University of California, Los Angeles, city=Los Angeles, postcode=90095, country=USA
[…]
[3]organization=School of Information and Communication Engineering, Communication University of China, city=Beijing, postcode=100024, country=China
[…]
[4]organization=College of Computer Science and Technology, Zhejiang University, postcode=310058, city=Hangzhou, country=China
[…]
[5]organization=School of Computing and Information Systems, Singapore Management University, postcode=178902, country=Singapore
[…]
[6]organization=School of Computing, National University of Singapore, postcode=119077, country=Singapore
@@
Temporal event forecasting aims to predict future events based on observed historical events  [1]. This is a fascinating task that paves the way for humans to master the operating rules of the world, meanwhile, it also possesses significant practical and application value, such as offering early warning to critical events like civil unrest or regional conflicts  [2, 3, 4]. Given the great values and impacts, temporal event forecasting has garnered increasing interest in both academic and research communities, and various studies have been conducted in recent years  [5, 6].
Figure 1: Illustration of leveraging LLM for temporal event forecasting. Given the complex event Israeli-Palestinian conflict , three formats of historical event representations, i.e., text (top side), graph (bottom side), or graph-text (both), can be fed into the LLMs, and the LLMs are expected to answer certain input questions about what will happend in the future.
@@
Among these studies, the representative formulations of temporal events can be broadly categorized into three formats [1]: graph-based, text-based, and graph-text hybrid. Specifically, the graph-based approach represents each event, i.e., the so-called atomic event  [7, 8, 9], in a structured format, i.e., a quadruple $(s,r,o,t)$, where $s$, $r$, $o$, and $t$ corresponds to the subject, relation  [^1] Relation refers to the type of the atomic event., object, and timestamp, respectively. As illustrated in Figure 1, each atomic event(e.g., (Palestine, use conventional military force, Israel, t-2)) is extracted from a textual document(e.g., Palestinian militant groups led by Hamas launch over 3,500 rockets from the Gaza Strip into Israel. ) , and multiple related atomic events form a complex event[10, 11, 6], such as the Israeli-Palestinian conflict. Following such a structured representation, which is also termed as Temporal Knowledge Graph (TKG), various methods [12, 13, 14, 15, 6] that target at modeling the temporal and relational patterns have been proposed and achieved remarkable progress [5]. In contrast, text-based methods, such as event script prediction [16] or forecast question answering (ForecastQA) [17], are characterized by directly consuming textual representations, which encapsulate more fine-grained details and contexts. Such detailed and contextual information is often ignored by graph-based methods since they are either not included in the ontology or not extracted by the information extraction system. Consequently, text-based event forecasting [17, 16] concentrates more on the capabilities of natural language understanding and reasoning. The third branch, i.e., graph-text hybrid method, aims to take advantage of the merits from both formats: structured reasoning and fine-grained information. Nevertheless, existing graph-text hybrid methods [4, 18, 11] primarily treat text as side information to be integrated into the graph-based backbones, without truly conducting reasoning and forecasting on the texts themselves.
@@
With the striking success of ChatGPT  [^2] https://chat.openai.com, LLMs have exhibited remarkable effectiveness across a wide range of tasks [19, 20, 21] as well as event forecasting [22, 23, 24, 25]. These work pioneer in applying LLMs to the task of event forecasting through in-context-learning (ICL) [23], instruction tuning [22, 25], and retrieval augmented generation (RAG) [26, 24] However, most of these approaches focus on discrete graph reasoning, thereby neglecting the broad contextual information inherent in natural language. Integrating the raw texts, from which the structured events are extracted, into the LLM-based forecasting processing is a natural and rational direction. Considering event forecasting is mostly related to critical scenarios such as international relationships and domestic stability, trustability and explainability are essentially required. Solely relying on the internal knowledge of LLMs would suffer from the problem of hallucination [23, 27]. While RAG [28, 29] presents a promising solution, current graph-based RAG methods [24, 26] are prone to noisy events obtained by unreliable event extraction systems. Therefore, the performance of text-based or graph-text hybrid RAG on temporal event forecasting is worth further exploration.
Furthermore, previous works [11] have identified that event forecasting suffers from severe popularity bias. Traditional graph-based methods, which rely on Graph Neural Networks (GNNs) or Temporal Logical Rules (TLR) to capture relational and temporal patterns from historical events, tend to prioritize popular entities. This bias becomes particularly problematic when predicting entities that are either appearing for the first time or are infrequently encountered [30], such as Egypt in Figure 1. In contrast, LLM-based methods have the potential to incorporate additional contextual information, potentially improving event forecasting for lesser-known entities. However, the extent to which LLMs, with their extensive world knowledge, are influenced by this bias remains uncertain.
@@
To address this gap, we propose a systematic investigation into the capabilities LLMs in temporal event forecasting. Nevertheless, the main obstacle lies in the lack of well-established benchmark datasets. Most existing works are using graph-only datasets, e.g., GDELT [2] and ICEWS [3], or text-only datasets, e.g., ForecastQA [17]. There are some text-enriched event graph datasets [11, 6], which enrich an existing structured TKG dataset with the raw texts that are downloaded through the news article URLs offered by GDELT. Unfortunately, the structured events in these datasets are highly noisy, because they are either from the original GDELT dataset [11] or extracted using poor-performing systems [6], with only about 50% accuracy as demonstrated in Section 1.2.
@@
In this work, to conduct the evaluation, we first construct an exploratory benchmark data based on a previous dataset MidEast-TE [6] and name our dataset as MidEast-TE-mini (short as MidEast-TE-m). Specifically, we sample a subset from the MidEast-TE [6] and extract structured events using the state-of-the-art (SOTA) LLM GPT-4  [^3] https://platform.openai.com/docs/models/gpt-4-and-gpt-4-turbo. Based on this dataset, we design a series of baseline methods to evaluate the performance of LLMs on temporal event forecasting. We are particularly interested in the following research questions: (1) How do LLMs perform across different formulations of event forecasting? (2) How does RAG perform on event forecasting? (3) How do LLMs perform regarding the popularity bias in event forecasting? To answer these research questions, we conduct extensive experiments to verify the functionality of every designed component. The experimental results unveil multiple interesting phenomena, which not only deepen our understanding of LLM-based event forecasting methods but also surface several interesting research directions. The main contributions of this work are as follows:
  • 1.
@@
    To the best of our knowledge, we are among the first to systematically evaluate the performance of LLMs in text-involved temporal event forecasting.
  • 2.
@@ 1.1 Problem Formulation
Conventional temporal event forecasting methods define each event as a quadruple $(s,r,o,t)$, referred to as an atomic event. Following recent studies [10, 6], atomic events are grouped into complex events (CE), which include a set of correlated atomic events and represent major events that span over longer time spans and cover more entities. For example, all the events within the red dashed box in Figure 1 illustrate a complex event, specifically the Israeli-Palestinian conflict. Consequently, we define an atomic event at timestamp $t$ as $(s,r,o,t,c)$, where $s\in\mathcal{E}$, $r\in\mathcal{R}$, and $o\in\mathcal{E}$ represent the subject, relation, and object, respectively; $\mathcal{E}$ and $\mathcal{R}$ denote the entity and relation sets, and $c$ represents the complex event type. Notably, the timestamp $t$ is converted into a relative timestamp. All atomic events at the same timestamp $t$ form an knowledge graph, denoted as $G_{t}={(s_{n},r_{n},o_{n},t,c_{n})}_{n=1}^{N}$, where $(s_{n},r_{n},o_{n},t,c_{n})$ represents the $n$-th event, and $N$ is the total number of events.
@@
In addition to the traditional graph-based representation, we aim to incorporate raw news text into temporal event forecasting. Specifically, each knowledge graph $G_{t}$ is extracted from a list of news documents $D_{t}=\left[d_{1},d_{2},\ldots,d_{k}\right]_{k=1}^{K}\in\mathcal{D}$, where $d_{k}$ denotes the $k$-th document at timestamp $t$ and $\mathcal{D}$ denotes the set of documents. It is important to note that the same document $d$ can contribute to multiple events. Given the historical knowledge graphs or news documents prior to $t$, denoted as $\mathbf{G}_{<t}=\left\{G_{0},G_{1},\ldots,G_{t-1}\right\}$ and $\mathbf{D}_{<t}=\left\{D_{0},D_{1},\ldots,D_{t-1}\right\}$, respectively, and a query, represented as $(s,r/o,t)$, temporal event forecasting aims to predict the missing object or relation.
1.2 Dataset Construction
@@
We first describe the data source that we utilized for curating the dataset. Then we present the dataset construction pipeline, followed by the dataset construction.
1.2.1 Data Source
@@
We construct our dataset based on a subset of MidEast-TE[6], therefore, we name our dataset as MidEast-TE-mini, short as MidEast-TE-m. The original MidEast-TE dataset includes both structured atomic events and news articles, which have already been clustered into complex events. Given the large scale of MidEast-TE, we sample a subset of complex events from it and keep the associated news articles as the original documents to build our dataset. Specifically, the time spans of the complex events in MidEast-TE are of high divergence, ranging from several days to more than three months. To reduce the potential bias introduced by outliers, we just sample 120 complex events whose time spans are within 40-60 days. For the shorter and longer complex events, we do not cover them in this exploratory dataset and leave them for future work.
Figure 2: The data distribution of MidEast-TE-m.
1.2.2 Construction Pipeline
@@
Event Extraction. We instruct GPT-4 to perform the task of event extraction. Since the set of event types is quite large ($>200$), putting all the event definitions into prompts yields high cost and may decrease the performance. Therefore, we follow the previous approach [6] and conduct a multi-level event extraction given the three-layer hierarchical structure of CAMEO [31] ontology. We include the description of each event type defined in CAMEO in the prompt. During the first-level extraction, we break down each article into multiple sentence-level trunks, each with around 150 tokens. Then, we extract the first-layer atomic events from each trunk. After recognizing atomic events based on first-level event types, for the second-level extraction, we group 15 atomic events with the same first-level event type and their respective trunks into a single prompt. This approach aims to maximally share the prompt while maintaining the high-quality extraction. The same procedure is then applied to the third-level extraction.
@@
1.2.3 Dataset Statistics
We split the dataset temporally and subsequently present its distribution. We then assess the quality of the benchmark dataset.
Data splits.
Table 1: Statistics of our curated dataset MidEast-TE-m.
@@ 1.2.3 Dataset Statistics
Data Distribution. Figure 2 (a) presents the distribution of the complex events, where each point represents a specific complex event. The x-axis indicates the number of atomic events contained within each complex event, while the y-axis represents the number of associated news documents. The size of each scatter point is proportional to the number of event entities within the complex event, and the color of the points corresponds to the time span of the events, with lighter colors representing shorter time spans and darker colors indicating longer durations. Apparently, more complex events typically encompass a larger number of atomic events and associated news texts. The variation in point size and color further highlights the diversity of complex events in terms of time span and the number of entities involved. In addition, the box plots above and to the right provide insights into the distribution of atomic events and news texts, respectively, within complex events. The Top-10 entities with the highest frequency in the dataset are shown in Figure 2 (b). In order to display the temporal trends of our dataset, Figure 2 (c) illustrates the temporal distribution of atomic events, news documents, and complex events across different months (x-axis) and years (y-axis). The color intensity reflects the number of events occurring in each time period. We can see that the temporal distribution of MIDEAST-TE-m is relatively uniform.
Table 2: Error analysis of the event extraction results in different datasets.
Dataset #atomic events Acc.(%) error type (%) time relation entity GDELT-TE 148 29.73 3.85 31.73 64.42 MidEast-TE 35 55.56 0 92.86 7.14 MidEast-TE-m(ours) 78 72.60 16.67 27.78 55.56
Human Evaluation. In order to evaluate the quality of the benchmark dataset, we conduct human evaluation. Specifically, given a news article and the extracted events, we ask the human evaluator to tell whether the extracted events are valid or not based on the original article.
We randomly sample 20 documents from the document set of the dataset, and conduct human evaluation based on the following criteria:
  • 1.
@@
To be noted, for every extracted atomic event, we evaluate it once by going through the three criteria sequentially. For example, if we identify the Time is incorrect, we stop the checking of the following two criteria. Table 2 shows the accuracy and error types of the event extraction results, including both our dataset MidEast-TE-m and previous datasets of MidEast-TE [6] and GDELT-TE [2]. To be noted, MidEast-TE is constructed using Vicuna-7b [32], while GDELT-TE is curated using proprietary extraction system. Clearly, the dataset constructed by our pipeline is the most accurate one. In our dataset constructed using GPT-4, there are the fewest extraction errors, which makes the following forecasting research more valid and reliable. However, we admit that an overall accuracy of 72.6% is still unsatisfactory in practice, and further efforts will be devoted to improved event extraction performance.
2 Methods